# The Illusion of Diminishing Returns: Measuring Long Horizon Execution in LLMs

### v1

@@ 3.1 Effect of increasing the number of turns
We first test our hypothesis that long-horizon execution can be challenging even on tasks where world-knowledge and planning are not required. We then study the benefits of scaling model size on long-horizon execution.
Setup. We evaluate the Qwen3 (Yang et al., 2025) and Gemma3 (Gemma-Team et al., 2025) model families, as they offer a range of sizes: [4, 8, 14, 32]B and [4, 12, 27]B parameters, respectively. For this experiment, we set the turn complexity to its simplest form ($K=1$), providing a single key per turn, and vary the number of turns. Models are instructed to output the final answer directly, without intermediate thinking tokens, with the format enforced via few-shot examples. In Appendix˜C we show the results below also hold with chain of thought and thinking models.
Result 1: Execution Alone is Challenging. We present the results in Figure˜4. All models except Gemma3-4B and Qwen3-4B achieve 100% accuracy on the first step, highlighting how they have the knowledge and reasoning capability required to perfectly do a single step of our task. Yet, task accuracy falls rapidly over subsequent turns. Even the best-performing model (Qwen3-32B) sees its accuracy fall below 50% within 15 turns. This confirms our hypothesis that long-horizon execution can be challenging for LLMs even when planning and knowledge requirements are removed.
Result 2: Non-Diminishing Benefit of Scaling Model Size. As shown in Figure˜4 (a), larger models sustain higher task accuracy for significantly more turns, resulting in a clear scaling trend for horizon length (Figure˜4 (c)). We abstain from deriving a “scaling law” since we can only obtain at most four model sizes from the same family, but the improvements do not seem diminishing. This observation is non-trivial. While the benefits of increasing model size are often attributed to improved knowledge capacity, our task is not knowledge-constrained, as even small models achieve perfect single-step accuracy, nor is the task more complex so that a larger model would be required. Yet, larger models are clearly more reliable at executing the task for longer. A possible explanation is the redundancy of internal circuits in larger models, which ensembles to reduce error (Lindsey et al., 2025). However, we find that simulating this redundancy with output-level aggregation of parallel compute (Appendix˜B) does not replicate the gains observed from scaling model size.
Takeaway 1. Long-horizon execution is challenging. Scaling model size significantly increases the number of turns a model can correctly execute.
@@ 3.2 Why Does Turn Accuracy Degrade? The Self-Conditioning Effect
One might expect a model’s per turn performance to remain constant. Yet, Figure˜4(b) shows the accuracy of individual turns steadily degrades as the number of turns increases. We investigate two competing hypotheses:
@@
Setup. To disentangle these factors, we conduct a counterfactual experiment by manipulating the model’s chat history. We control the error rate by injecting artificial output histories with a chosen error rate in the same format. If we fully heal the history, with a $0\%$ error rate, degradation in the model’s turn accuracy between turn 1 and a later turn can be attributed to long-context issues. If a model’s accuracy for a fixed later turn consistently worsens with increasing error rate in prior turns, this would demonstrate that models condition on their past mistakes, increasing the likelihood of future errors.
@@
Result 4: Unlike long-context, scaling model size does not mitigate self-conditioning. Notice that the accuracy at turn 100 at the induced error rate of 0 consistently improves for larger models. As shown in Figure˜5 (c), scaling to frontier (200B+ parameter) models like Kimi-K2 (Kimi-Team et al., 2025), DeepSeek-V3 (DeepSeek-AI et al., 2025), and Qwen3-235B-Instruct-2507 (Yang et al., 2025) largely solves long-context degradation for up to 100 turns, achieving near-perfect accuracy on a healed history. However, even these large models remain susceptible to self-conditioning, as their performance consistently degrades as the induced error rate in their history increases. This may be akin to recent results showing larger models shift in personality during multi-turn conversations (Choi et al., 2024; Becker et al., 2025), where in our case, the drift is toward a personality that makes errors.
Figure 5: Language models self-condition on their previous mistakes, leading to more mistakes in subsequent turns. By manipulating the chat history, we counterfactually vary the fraction of errors in previous turns. We find this increases the likelihood of errors in future turns (left). This shows a source of degradation in turn-wise model accuracy beyond long-context, as in the turn 100 slice (right) model accuracies are much higher when we provide a fully correct history. Scaling model size increases self-conditioning, even for frontier non-thinking models.
Takeaway 2. Models self-condition on their previous mistakes, leading to degradation in per-step accuracy. Scaling model size is not sufficient to mitigate this.
@@
We now study the effect of enabling sequential test time compute (“thinking”) for these models.
Setup. We enable thinking for the Qwen3 models, which are post-trained with reinforcement learning (RL). These models are trained to generate reasoning traces even when the context contains only the final answers from previous turns. This contrasts with standard chain-of-thought prompting, where models often fail to reason if prior reasoning steps are omitted from the context. We found that the Gemma3 models, when prompted for CoT, were unable to generate reasoning if prior traces were omitted from the context. These models exhibited a form of format-based self-conditioning: after observing a history of turns with only final outputs, they would ignore explicit user instructions to think step by step and revert to producing only a final answer, further discussed in  Appendix˜G. Given this limitation of CoT prompted Gemma3 models, we just focus on studying the self-condition effect on thinking-enabled Qwen3 models, by observing turn 100 accuracy while controlling the error rate in prior turns as before.
Result 5: Thinking Fixes Self-Conditioning. In Figure˜6, we clearly find that the Qwen3 thinking models do not self-condition—the models’ accuracy at turn 100 remains stable, regardless of the error rate in its context. This could arise from two reasons. First, RL training can reduce the most likely next token prediction behaviour of language models, making them oriented towards task success rather than continuing the context. Second, the removal of thinking traces from prior turns could reduce the influence of prior turns on the model’s output, as it thinks about the new turn independently. By inspecting the models’ thinking traces, we observe that they do not refer back to prior turns in their chain of thought. Furthermore, we experiment with context management by explicitly removing prior history as a potential fix, and find that it indeed mitigates self-conditioning (Appendix˜A).
@@ 3.3 What is the length of tasks models can complete in a single turn?
In the previous sections, we measured how many turns models can successfully execute a single retrieve-then-compose step. However, most real-world tasks require more complex processing every turn. The total task length a model can handle is a function of both the number of turns and the number of steps to execute per turn. We now measure the latter dimension: the maximum number of steps a model can execute per turn.
@@
Figure 7: Benchmarking the length of task models can execute in a single turn. Without CoT or thinking, even the biggest models to fail to execute more than a few steps (left). Sequential test time compute (thinking tokens) significantly improve this, especially when trained with RL (eg R1 vs DeepSeek V3) (right). GPT-5 is far ahead of the rest, executing over 1000 steps, with Claude-4-Sonnet second at around 400.
@@
Setup. To quantify this capability, we propose a benchmark that measures the maximum single-turn execution capacity of various models. We run a binary search (Lehmer, 1960) to find the highest turn complexity ($K$, the number of keys) the model can provide the correct sum for with accuracy $\geq 80\%$. We evaluate a suite of frontier models like GPT-5 (OpenAI, 2025), Claude-4 Sonnet (Anthropic, 2025), Grok 4 (xAI, 2025), Gemini 2.5 Pro (Gemini Team, 2025), Kimi K2 (Kimi-Team et al., 2025), Qwen3-Instruct-235B-2507 (Yang et al., 2025), and DeepSeek R1 (Guo et al., 2025). An advantage of our benchmark is that it is contamination-free, as new examples can be generated programmatically.
@@
Result 1: Without chain of thought, non-thinking models struggle to chain even two steps in a single turn. In Figure˜12 (left), we first find that when prompted to directly answer, without chain-of-thought, the larger Qwen3 32B, Gemma3 27B, as well as frontier non-think models like DeepSeek-V3 (670B), and Kimi K2 (1026B), fail to execute even a turn complexity of 2. This is consistent with prior work showing the necessity of thinking tokens for transformers to perform sequential tasks (Weiss et al., 2021; Merrill and Sabharwal, 2023). We highlight this because many agentic workflows directly ask the model to act, without chain-of-thought, to fit more actions in the context window. We see that the number of steps the model can execute in a single turn improves significantly with chain-of-thought. It shows the importance of reasoning before acting (ReAct (Yao et al., 2023)) for agents. In Appendix˜B, we also show that parallel test time compute like majority voting (Wang et al., 2022) only leads to marginal improvements in both single turn execution length and number of turns. This provides preliminary evidence that for long-horizon execution, sequential test-time compute is more effective.
@@
Result 2: Benchmarking Frontier Models. In Figure˜12 (right), we benchmark frontier models on the length of task they can execute in a single turn. We find a surprisingly large gap between GPT-5 (codenamed Horizon) and others like Gemini 2.5 Pro, Grok 4, and DeepSeek R1. We also find that the RL-trained thinking model DeepSeek R1 substantially outperforms its instruction-tuned counterpart, DeepSeek-V3. Overall, long-horizon execution is a challenge in which open-weight models are still catching up to those only available through APIs, highlighting opportunities for future research.
Takeaway 3. Thinking models fix self-conditioning, and can also execute significantly longer tasks in a single turn.
@@ 4 Related Work
Long Context. Much prior work has focused on improving the maximum context length that can be provided in the input to a language model (Su et al., 2021), and evaluating whether (Tay et al., 2020) and how (Olsson et al., 2022; Li et al., 2023) models maintain performance as the context gets longer (Tay et al., 2020). Closest is the recent RULER (Hsieh et al., 2024) and GSM-Infinite (Zhou et al., 2025b), which also uses synthetic data to systematically evaluate long-context abilities. While long-context will help models execute for longer, it is a different capability compared to long-horizon execution (Zhou et al., 2023; Chen et al., 2024a), as it focuses on performance as a function of input, not output length. We identified one such difference, the self-conditioning effect–where past errors in model output increase the chance of future mistakes, and disentangle this effect from long-context degradation in  Section˜3.2.
Classical Reasoning and Planning. Automated planning, especially when formulated for discrete and deterministic spaces, has long been a mainstay of artificial intelligence research. In symbolic AI, once tasks are formalized, for example into STRIPS plans (Fikes and Nilsson, 1971), they can be evaluated without issues in execution. Prior work (Chen et al., 2024b; Valmeekam et al., 2024) has shown LLMs struggle to match symbolic algorithms for automated planning. In contrast, we focus on straightforward execution of provided plans over a long horizon.
@@
Increasing Task Complexity (length). Multiple works have recently shown how models worsen as “problem complexity” increases Zhou et al. (2025b), often attributed to failures of reasoning (Cheng, 2025; Shojaee et al., 2025). Recently, multiple real-world long-horizon agentic benchmarks have been proposed (Backlund and Petersson, 2025; Xie et al., 2024; Shen et al., 2025), where prior work has studied planning failures (Chen et al., 2024b). By designing a task where no reasoning is required, given that we provide the model the requisite plan and knowledge, we show that execution alone can be a challenge (Zhu et al., 2025; Sun et al., 2025), degrading model accuracy on longer tasks. Our observations on scaling could hold for the related problem of length-generalization—training models to succeed on tasks longer than those seen during training (Fan et al., 2024; Cai et al., 2025).
Controlled Evaluations with Synthetic Data. Improvements on real-world benchmarks are the ultimate measure of AI progress, but understanding LLM capabilities and shortcomings sometimes requires disentangling the many factors that compound in real tasks. Our empirical approach of performing a controlled study of LLM capabilities, using a simplified task to remove confounders, aligns with recent work on architecture design (Allen-Zhu, 2024; Poli et al., 2024), recall from parametric memory (Arora et al., 2023) (where our retrieve step requires in-context retrieval), length generalization (Lee et al., 2025), and the ability to form new abstractions (Chollet et al., 2024). We focus on a different capability—long-horizon execution—which we posit is becoming increasingly important as we enter the era of experience (Silver and Sutton, 2025).
5 Discussion
@@
Scaling laws for language models show diminishing returns on the loss for the single step of predicting the next token (Kaplan et al., 2020; Hoffmann et al., 2022). When models competed in simple knowledge-based question-answering tasks such as MMLU (Hendrycks et al., 2020), such single-step measurements could inform us about the rate of progress. This has changed in the last year. Where earlier we could only post-train on human demonstrations (Mishra et al., 2021), language models can now be trained with just rewards (Shao et al., 2024), enabling sophisticated reasoning (Guo et al., 2025; Jain et al., 2024) and agents (Kimi Team et al., 2025). This opens up the opportunity to solve much longer tasks where earlier human supervision would be too expensive to scale. Our work shows how diminishing returns on single-step performance can compound to provide large benefits in the length of tasks a model can solve. This motivates the need to study empirical scaling laws for horizon length in agents (Hilton et al., 2023). An astute reader might wonder if execution failures should be solved by providing the model access to tools (Schick et al., 2023). Tools indeed help shift the burden of execution from probabilistic models to reliable programs. However, reasoning is often fuzzy, and not always easy to implement as a tool, requiring the model to execute some by itself. Even calling the right tools requires reliable execution from the model (Patil et al., 2025).
@@
While there has been recent interest in evaluations for LLM reliability (Vendrow et al., 2025; Yao et al., 2024), they do not focus on long-horizon outputs, where the context differs every step. For example,  Yao et al. (2024) focus on the $pass^{k}$ metric, which checks if the model makes a mistake when we sample $k$ generations. However, this keeps the input fixed, and at 0 temperature (deterministic sampling), becomes equivalent to $pass@1$. In long-horizon tasks, error compounds irrespective of temperature, as shown in  Figure˜13. Further, we find that even single-step error rates can grow as the output length increases. While this might seem similar to prior work showing degradation in long-context (Zhou et al., 2025a), using counterfactual experiments, we show this is rather due to the model self-conditioning on its own generated mistakes.

### v2

@@ 3.1 Effect of increasing the number of turns
We first test our hypothesis that long-horizon execution can be challenging even when a model has the required knowledge and planning ability, and then study the benefits of scaling model size.
Setup. We evaluate the Qwen3 (Yang et al., 2025) and Gemma3 (Gemma-Team et al., 2025) model families, as they offer a range of sizes: [4, 8, 14, 32]B and [4, 12, 27]B parameters, respectively. For this experiment, we set the turn complexity to its simplest form ($K=1$), providing a single key per turn, and vary the number of turns. Models are instructed to output the final answer directly, without intermediate thinking tokens, with the format enforced via few-shot examples. We verify that format-following errors are not the primary failure mode (Appendix˜F). We also show that the results below hold with chain-of-thought (CoT) prompting, and thinking models (Appendix Figure˜12), and the trends are not affected by the temperature used (Appendix Figure˜13).
Result 1: Execution alone is challenging. As seen in Figure˜4 (a), all models except Gemma3-4B and Qwen3-4B achieve near-perfect accuracy on the first step, confirming they have the knowledge required to perfectly do a single step of our task. Yet, task accuracy falls rapidly over subsequent turns (Figure˜4 (c)). Even the best-performing model (Qwen3-32B) sees its accuracy fall below 50% within 15 turns. This confirms our hypothesis that long-horizon execution can be challenging for LLMs, even if they have the needed knowledge and plan.
Result 2: Non-diminishing benefits of scaling model size. As shown in Figure˜4 (c), larger models sustain higher task accuracy for significantly more turns, resulting in a clear scaling trend for horizon length (Figure˜4 (b)). We abstain from deriving a “scaling law” since we can only obtain at most four model sizes from the same family, but the improvements do not seem diminishing. This observation is non-trivial. While the benefits of increasing model size are often attributed to improved capacity for knowledge, our task is not knowledge-constrained, as models achieve near-perfect first step accuracy (Figure˜4 (a)), nor is the task more complex so that a larger model would be required. Yet, larger models are clearly more reliable at executing the task for longer. A possible explanation is the redundancy of internal circuits in larger models, which ensembles to reduce error (Lindsey et al., 2025). However, we find that simulating this redundancy with output-level aggregation of parallel compute does not replicate the gains observed from scaling model size (Appendix˜B).
Figure 4: Scaling model size has non-diminishing improvements in the number of turns it can execute. The first-step accuracy for our task is near-perfect for all except the smallest models (a). Yet, as the model size is scaled, the horizon length increases significantly (b). We also see the effect of scaling in widening the gap between small and large models in task accuracy (c) and turn accuracy (d) as the number of turns increases. The shaded region is the mean $\pm$ one standard deviation over 100 samples; the solid line is the moving average over 5 turns; the dotted line is a hypothetical baseline model with constant step-accuracy of 0.99.
@@ 3.2 Why Does Turn Accuracy Degrade? The Self-Conditioning Effect
One might expect a model’s turn accuracy to remain constant. Yet, Figure˜4 (d) shows the accuracy of individual turns degrading as the number of turns increases. We investigate two competing hypotheses:
@@
Setup. To disentangle these factors, we conduct a counterfactual experiment by manipulating the model’s chat history. We control the error rate by injecting artificial output histories with a chosen error rate in the same format. If we fully heal the history, with a $0\%$ error rate, degradation in the model’s turn accuracy between turn 1 and a later turn can be attributed to long-context issues. If a model’s accuracy for a fixed later turn consistently worsens with increasing error rate in prior turns, this would support our self-conditioning hypothesis.
@@
Result 4: Unlike long-context, scaling model size does not mitigate self-conditioning. At the error rate of 0%, notice that the accuracy at turn 100 consistently improves as you scale model size. As shown in Figure˜5 (b), scaling to frontier (200B+ parameter) models like Kimi-K2 (Kimi-Team et al., 2025), DeepSeek-V3 (DeepSeek-AI et al., 2025), and Qwen3-235B-Instruct-2507 (Yang et al., 2025) largely solves long-context degradation for up to 100 turns, achieving near-perfect accuracy on a healed history. However, even these large models remain susceptible to self-conditioning, as their performance consistently degrades as the induced error rate in their history increases. This may be akin to recent results showing larger models shift in personality during multi-turn conversations (Choi et al., 2024; Becker et al., 2025), where in our case, the drift is toward a personality that makes errors.
Figure 5: Models self-condition on their previous mistakes, leading to more mistakes in subsequent turns. By manipulating the chat history, we counterfactually vary the fraction of errors in previous turns. We find this increases the likelihood of errors in future turns (left). This shows a source of degradation in turn-wise model accuracy beyond long-context, as in the turn 100 slice (right) model accuracies are much higher when we provide a fully correct history. Scaling model size increases self-conditioning, even for frontier non-thinking models.
In Appendix˜G, we try the above setup of output manipulations with CoT prompting, finding that accuracy still deteriorates as the induced error rate increases. A potential confounder is that the manipulated outputs deviate from the CoT. We try to mitigate this issue with programmatically generated CoT traces, but still observe self-conditioning. We also try removing CoT traces from previous turns from history, which causes the model to stop using CoT completely. So, we now present results for the Qwen3 thinking models, which are trained with reinforcement learning (RL) to think even when previous turn traces are not presented. As before, we observe their turn 100 accuracy while controlling the error rate in prior turns.
@@
Result 5: Thinking fixes self-conditioning. In Figure˜6, we observe that the Qwen3 thinking models do not self-condition—the accuracy of the models at turn 100 remains stable, regardless of the error rate in its context. This could arise from two reasons. First, RL training can reduce the most likely next token prediction behaviour of language models, making them oriented towards task success rather than continuing the context. Second, the removal of thinking traces from prior turns could reduce the influence of prior turns on the model’s output, as it thinks about the new turn independently. By inspecting the models’ thinking traces, we observe that they do not refer back to their answers in prior turns, which could be a potential reason why they do not self-condition. Inspired by this, for non-thinking models, we experiment with context engineering by explicitly removing prior history and find that it indeed mitigates self-conditioning (Section˜A.2). We also find that just prompting models to self-verify their answers does not solve self-conditioning completely (Section˜A.1).
@@ 3.3 What is the length of tasks models can complete in a single turn?
In the previous sections, we measured how many turns models can successfully execute a single retrieve-then-compose step. However, most real-world tasks require more complex processing every turn. In fact, in Appendix˜C we show that the horizon length of different models can vary significantly at different turn complexities. The total task length a model can handle is a function of both the number of turns and the number of steps to execute per turn. We now measure the latter dimension: the maximum number of steps a model can execute per turn.
@@
Figure 7: Benchmarking the length of task models can execute in a single turn. Without CoT or thinking, even the biggest models fail to execute more than a few steps (a). Sequential test time compute (thinking tokens) significantly improves this, especially when trained with RL (b), where GPT-5 is far ahead of the rest.
@@
Setup. To quantify the length of task models can complete in one go, without user input, we run a binary search (Lehmer, 1960) to find the highest turn complexity ($K$, the number of keys) the model can provide the correct sum for with accuracy $\geq 80\%$. We evaluate a suite of frontier models like GPT-5 (OpenAI, 2025), Claude-4 Sonnet (Anthropic, 2025), Grok 4 (xAI, 2025), Gemini 2.5 Pro (Gemini Team, 2025), Kimi K2 (Kimi-Team et al., 2025), and DeepSeek-R1 (Guo et al., 2025). An advantage of our benchmark is that it is contamination-free, as new examples can be generated programmatically.
@@
Result 6: Without CoT, non-thinking models struggle to chain more than a few steps per turn. In Figure˜7 (left), we first find that when prompted to answer directly, without chain-of-thought, the larger Qwen3 32B, Gemma3 27B, as well as frontier non-thinking models like DeepSeek-V3 (670B), and Kimi K2 (1026B), fail to execute even a turn complexity of more than six. This is consistent with prior work showing the necessity of thinking tokens for transformers to perform sequential tasks (Weiss et al., 2021; Merrill and Sabharwal, 2023). We see that the number of steps the model can execute in a single turn improves significantly with chain-of-thought. This reinforces the importance of reasoning before acting (ReAct (Yao et al., 2023)) for agents, even if this costs more and fills up the context window. We show preliminary evidence that parallel test time compute is not as helpful, with majority voting leading to only marginal improvements in execution length (Appendix˜B).
@@
Result 7: Benchmarking frontier models. In Figure˜7 (right), we benchmark frontier models on the length of task they can execute in a single turn. We find a surprisingly large gap between GPT-5 (codenamed Horizon) with 2176 steps and others like Claude-4 Sonnet (432 steps), Grok 4 (384 steps), and Gemini 2.5 Pro (120 steps). Overall, even our simple task can separate frontier models in their long-horizon execution capability, and presents a clear opportunity to improve current open-weight models.
@@ 4 Related Work
Increasing Task Length. Multiple works have recently shown how models worsen as problem complexity increases (Zhou et al., 2025b), often attributed to failures of reasoning (Cheng, 2025; Shojaee et al., 2025). Recently, multiple real-world long-horizon agentic benchmarks have been proposed (Backlund and Petersson, 2025; Xie et al., 2024; Shen et al., 2025), where prior work has studied planning failures (Chen et al., 2024b). By designing a task where no reasoning is required, given that we provide the model the requisite plan and knowledge, we show that execution alone can be a challenge, degrading model accuracy on longer tasks. Our observations on scaling could hold for the related problem of length-generalization—training models to succeed on tasks longer than those seen during training (Fan et al., 2024; Cai et al., 2025).
Long Context. Much of prior work has focused on improving the maximum context length that can be provided in the input to a language model (Su et al., 2021), and evaluating whether (Tay et al., 2020) and how (Olsson et al., 2022; Li et al., 2023) models maintain performance as the context gets longer (Tay et al., 2020). Closest is the recent RULER (Hsieh et al., 2024) and GSM-Infinite (Zhou et al., 2025b), which also use synthetic data to systematically evaluate long-context abilities. While long-context will help models execute for longer, it is a different capability compared to long-horizon execution (Zhou et al., 2023; Chen et al., 2024a), as it focuses on performance as a function of input, not output length. We identified one such difference, the self-conditioning effect—where past errors in model output increase the chance of future mistakes, and disentangle this effect from long-context degradation in Section˜3.2. Motivated by a similar effect at the token level, termed exposure bias, Ranzato et al. (2015) proposed training language models with RL.
@@
Scaling LLMs and RL. Scaling laws for language models show diminishing returns on the loss for the single step of predicting the next token (Kaplan et al., 2020; Hoffmann et al., 2022). When models competed in simple knowledge-based question-answering tasks such as MMLU (Hendrycks et al., 2020), such single-step measurements could inform us about the rate of progress. This has changed in the last year. Where earlier we could only post-train on human demonstrations (Mishra et al., 2021), language models can now be trained with just rewards (Shao et al., 2024), enabling sophisticated reasoning (Guo et al., 2025; Jain et al., 2024) and agents (Kimi Team et al., 2025). This opens up the opportunity to solve much longer tasks where earlier human supervision would be too expensive to scale. Our work shows how diminishing returns on single-step performance can compound to provide large benefits in the length of tasks a model can solve. This motivates the need to study empirical scaling laws for horizon length in agents (Hilton et al., 2023).
Tool Use. In symbolic AI, once tasks are formalized, for example, into STRIPS plans (Fikes and Nilsson, 1971), they can be executed without issues. Prior work (Chen et al., 2024b; Valmeekam et al., 2024) has shown LLMs struggle to match symbolic algorithms for automated planning. In contrast, we show LLMs can fail on straightforward execution (Zhu et al., 2025; Sun et al., 2025; Stechly et al., 2024) over a long horizon even when the plan is provided. Teaching LLMs to use tools offers one way to shift the burden of execution from probabilistic models to reliable programs (Schick et al., 2023). However, reasoning is often fuzzy and not always easy to implement as a tool, requiring the model to execute some steps by itself. Even calling the right tools requires reliable execution from the model (Patil et al., 2025).
@@
Reliability. While there has been recent interest in evaluations for LLM reliability (Vendrow et al., 2025; Yao et al., 2024), it does not focus on long-horizon outputs, where the context differs every step. For example,  Yao et al. (2024) focus on the $pass^{k}$ metric, which checks if the model makes a mistake when we sample $k$ generations. However, this keeps the input fixed, and at 0 temperature (deterministic sampling), becomes equivalent to $pass@1$. In long-horizon tasks, error compounds irrespective of temperature, as shown in Appendix Figure˜13.
5 Conclusion
In this work, we show how short-task benchmarks may give the illusion of slowing progress for modern language models. We show that scaling model size increases the number of turns a model can execute, while sequential test-time compute increases the length of tasks a model can perform on a single turn. Together, these contribute to dramatically increasing horizon lengths for LLMs.